# Topological Structure Learning Should Be A Research Priority for LLM-Based Multi-Agent Systems

### v1

@@ Abstract
Large Language Model-based Multi-Agent Systems (MASs) have emerged as a powerful paradigm for tackling complex tasks through collaborative intelligence. Nevertheless, the question of how agents should be structurally organized for optimal cooperation remains largely unexplored. In this position paper, we aim to gently redirect the focus of the MAS research community toward this critical dimension: develop topology-aware MASs for specific tasks. Specifically, the system consists of three core components—agents, communication links, and communication patterns—that collectively shape its coordination performance and efficiency. To this end, we introduce a systematic, three-stage framework: agent selection, structure profiling, and topology synthesis. Each stage would trigger new research opportunities in areas such as language models, reinforcement learning, graph learning, and generative modeling; together, they could unleash the full potential of MASs in complicated real-world applications. Then, we discuss the potential challenges and opportunities in the evaluation of multiple systems. We hope our perspective and framework can offer critical new insights in the era of agentic AI.
@@ 1 Introduction
Large Language Model (LLM) agents [1, 2, 3] have demonstrated strong capabilities across diverse tasks, such as code generation [4, 5, 6], mathematical reasoning [7, 8], and social simulation [9, 10]. Building on the impressive capabilities of single-agents in general tasks, LLM-based Multi-Agent Systems (MASs) [11] have emerged as a powerful paradigm that coordinate multiple specialized agents to collaboratively solve more complex tasks. These systems assign distinct roles—such as planners, retrievers, and reasoners–to distinct agents, enabling human-like collaboration through a division of labor. Following the analogy of optimizing a human organization, an agentic system can be seen as a dynamically evolving ensemble that seeks a moving equilibrium: whenever new agents join the system or existing ones depart, the ensemble reconfigures itself to maximize its long-term utility subject to budget and latency constraints.
To enable inter-agent communication, coordination, and role specialization, MAS architectures must support agents that are not only individually competent but also capable of collaborating through well-orchestrated interaction patterns. Despite the promising progress made in the field, the question of how agents should be organized to support effective and efficient cooperation remains largely underexplored [12]. This reveals an urgent need to rethink the multi-agent design: What topological structures of LLM-based MASs are best to accomplish an unseen, complex task?
Why topology matters? In MASs, effective collaboration depends not only on the capabilities of individual agents but also on how they are organized and coordinated. Typically, as task complexity increases, larger groups of agents are required, leading to a significant rise in communication overhead. As revealed by recent studies, naïve scaling without adaptive structural design often results in redundant communication and suboptimal coordination [12, 13, 14].
(a) GPT-3.5
(b) GPT-4o
@@
Figure 1 shows that task performance varies by up to 10% if using different topologies. Furthermore, the choice of communication structure carries critical safety implications, whereby poorly designed topology may lead to misinformation propagation, biased reasoning, or the generation of harmful content by agents [15].
@@
What are the key challenges? First, MASs are highly task-dependent. Reasonable topological structures often depend on the characteristics of a specific task; yet, these characteristics are often implicit, difficult to model and to generalize, thereby hindering the development of universal topology designs. Some prior works attempt to explore dynamic topologies, but remain in the early stages and still lack the ability to adapt topological structures based on queries or task demands [16, 17]. Second, the design space of topological structures is inherently combinatorial, involving both agent selection and inter-agent connectivity, which grows exponentially with increasing agents. This makes efficient structure optimization computationally prohibitive as the system scales. Structure pruning techniques like edge pruning [13] and agent pruning [14, 17] are often applied in a static and one-off manner, lacking the ability to model the dynamic interdependencies and interactions among agents. Third, unlike in single-agent settings, the effectiveness of a given topological structure in MASs cannot be evaluated in isolation; rather, its performance must be evaluated through complete system execution within a specific task context, which entails substantial runtime overhead.
What should be learned and how? We summarize three core components in MASs: agents, communication links, and the overall topology that determines how agents are connected and interact. The candidate subset of agents should be capable and complementary agents for a given task, while communication links should be optimized to eliminate redundant message passing and minimize communication overhead. The overall topology should align with the coordination pattern of tasks to enable efficient information flow. In this position paper, we advocate a structure-aware paradigm that treats topological design as a key element of LLM-based MASs. We propose a unified framework (Figure 2) that decomposes topological structure optimization into three stages: agent selection, structure profiling, and topology synthesis, each corresponding to a core aspect of the topological structure. From this perspective, we reveal critical limitations of existing approaches and outline future directions toward more adaptive and efficient multi-agent architectures.
@@
Potentially improved applications. Current agent-based LLM applications can potentially benefit from a topology-aware structure. For instance, in customer support automation, dynamic agent topologies can reduce response latency, minimize redundant information routing, and improve relevance in multi-turn dialogues for support systems involving billing, troubleshooting, and multilingual hand-offs. For collaborative code generation, task-specific topologies can support better coordination among agents with distinct roles (planner, coder, tester), which adapt as tasks evolve. Moreover, in scientific discovery, research assistants powered by LLMs can benefit from topologies that optimize how domain-specific agents (e.g., literature retrievers, experimental planners, hypothesis evaluators) interact, particularly when solving open-ended scientific queries. Given the dynamic and flexible nature of interaction topology, more applications will emerge and benefit.
@@
The organization of this paper is as follows: we first introduce the basics of LLM-based MASs in Section 2. We then articulate the proposed research framework for topology optimization with open problems and research directions in Section 3. Subsequently, Section 4 presents challenges and opportunities in evaluation, followed by discussions on multiple systems in Section 5. Finally, Section 6 concludes the paper. Note that due to space limitations, more details of our proposed research directions are in the appendix.
2 LLM-based Multi-Agent Systems
While LLM-based single-agent systems have obtained notable success on a wide range of tasks, complicated problems can exceed the capabilities of a single model. This motivates the development of MASs, where multiple LLM-based agents collaborate through natural language communication to constructively divide roles, specialize functions, and jointly solve tasks. To support such collaboration, we define an LLM-based MAS as a set of autonomous agents instantiated from LLMs, denoted as $\hat{\mathcal{A}}=\{a_{1},...,a_{i}\}\subseteq\mathcal{A}$. Each agent $a_{i}\in\mathcal{A}$ is assigned a specific role, such as planner, coder, or verifier, and interacts with other agents to collectively solve tasks. Appendix A shows our notations.
In practice, MASs adopt various structural topologies, such as chains, stars, trees, or densely connected graphs, each encoding different reasoning workflows and communication costs. For instance, chain structures enable step-by-step refinement (e.g., MetaGPT [18]), while tree structures support hierarchical planning (e.g., AutoGen [1]). Star topologies simplify control via centralized agents (e.g., CAMEL [2]), and fully connected graphs maximize information sharing (e.g., GPTSwarm [19]). Each structure presents distinct trade-offs in coordination efficiency, scalability, and reasoning capability, and is inherently better suited to different tasks depending on its specific communication patterns and collaborative requirements. Formally, we model the collaboration structure as a directed graph $\mathcal{G}=(\mathcal{A},\mathcal{E})$, where $\mathcal{E}\subseteq\mathcal{A}\times\mathcal{A}$ represents the communication links. At each round $t$, agent $a_{i}$ receives messages from its in-neighbors and generates an updated response $R_{i}^{(t)}=a_{i}(P_{\text{sys}},P_{\text{user}}^{(t)})$, where $P_{\textit{sys}}$, $P_{\text{user}}$ denote the system prompt and user prompt at round $t$ respectively. $P_{\text{user}}^{(t)}=P_{\text{user}}^{(0)}\cup\mathcal{O}_{i}^{(t-1)}$, where $\mathcal{O}_{i}^{(t-1)}$ denotes the aggregated messages from neighbors of agent $a_{i}$.
@@ 2 LLM-based Multi-Agent Systems
Figure 2: The proposed framework: agent selection, structure profiling, and topology synthesis.
3 Holistic Research Framework
@@
In this section, we propose a framework to learn the optimal topological structure for MAS. Our framework consists of three sequential stages: agent selection, structure profiling, and topology synthesis, as illustrated in Figure 2. Each stage has the potential to trigger new open problems and research directions. Specifically, ❶ Given a new task, a pool of candidate agents with diverse capabilities is instantiated and made available. The first step is to select a subset of agents $\hat{\mathcal{A}}\subseteq\mathcal{A}$ best suited to collaborate on the task, based on factors such as skill specialization, role diversity, and past performance. This determines “who” participates in the communication topology. ❷ For each task, the system identifies a macro structure as a communication pattern, which captures abstract interaction forms, such as chains, stars, or trees. These macro structures reflect coarse-grained collaboration strategies and serve as an initial prototype for the following topology synthesis. ❸ Based on steps ❶ and ❷, the final stage synthesizes a concrete and directed communication graph, which we reference as the micro structure. It specifies the exact flow of information among the selected agents under the given macro structure, and determines which agents exchange messages and in what direction.
End-to-end Optimization. While we argue that the entire problem should be decomposed into three stages due to its challenging nature, it does not mean that end-to-end optimization is impossible. These three stages are tightly coupled and can be optimized jointly, adopting an “Expectation-Maximization” paradigm [20]. For instance, the agent selection algorithms might not be optimal, which can then be refined by the following procedures. On the other hand, the feedback from structure profiling could also have some impact on the micro-level structure of the selected agents. The flexibility of an end-to-end algorithm enables the system to dynamically refine all components in a closed-loop manner, adapting to changing task requirements and coordination dynamics.
Due to the high cost and complex applications for MASs, however, it is easier to start with a “divide-and-conquer” manner to decompose it into the above three stages. In the following, we articulate how to learn each stage separately. The decomposition allows us to provide deeper insights into each aspect, while the overall framework remains flexible enough to support joint optimization across stages.
@@ 3.1 Agent Selection
Selecting proper agents is a cost-utility optimization problem that must balance exploration (testing alternatives) and exploitation (using the best-known option).
Users must decide which agents to choose as candidates: high-performance but expensive SaaS LLMs versus smaller local checkpoints, and how many queries to assign to each agent. Under-exploration risks missing superior models, while over-exploration and exhaustive trials are extremely expensive.
@@
Existing agent selection schemes are fundamentally limited by three major challenges: Single-Agent Focus, Structural Myopia, and Lack of Scalability.
Users must balance when generalists and specialists fit into a particular task pipeline.
Position. We advocate a structure-aware, lifecycle-aware approach to agent selection that optimizes for multi-objective reward (task success, cost, latency, safety), which should remain adaptive to catalog churn and evolving task distributions.
We propose two viable solutions, Static Selection and Dynamic Selection, depending on the specific scenarios. When task distribution, communication structure, evaluation metrics, costs, and the LLM catalog remain unchanged over the deployment horizon, e.g.,  nightly batch summarization with a locked-in set of SaaS APIs, where the utility of an agent can be seen as stationary, the problem reduces to estimating and selecting an optimal subset based on the interaction structure, and thus Static Selection is preferred. Formally, assuming that each agent in the pool is associated with a cost $C(a_{i})\in\mathbb{R}_{\geq 0}$ (e.g., API credits, GPU runtime), for a downstream task $x$ under macro structure $M$, we can define a utility function $\mathcal{U}:(x,M,\hat{\mathcal{A}})\longmapsto[0,1]$ to evaluate the collective performance of a subset $\hat{\mathcal{A}}\subseteq\mathcal{A}$ on task $x$. This utility can reflect metrics such as answer accuracy, semantic relevance, or user engagement (e.g., click-through rate). The goal of static selection to find $k$ agents to balance performance-cost trade-offs:
$\mathcal{A}^{\star}=\underset{\hat{\mathcal{A}}\subseteq\mathcal{A},|\hat{\mathcal{A}}|=k}{\mathrm{argmax}}\Bigl{[}\mathbb{E}_{x\sim\mathcal{D}}\bigl{[}\mathcal{U}(x,M,\hat{\mathcal{A}})\bigr{]}-\sum_{a_{i}\in\hat{\mathcal{A}}}\gamma(a_{i})C(a_{i})\Bigr{]},$ (1)
@@
The multiplier $\gamma$ in Equation 1 can be treated as a stochastic, context-dependent signal, enabling adaptive selection policies that dynamically balance performance and economy as conditions drift.
Multi-armed bandits (MAB) and their combinatorial variants offer a principled framework to efficiently explore promising agents and exploit high-utility subsets. Unlike reinforcement-learning (RL) training loops that continuously update a policy’s parameters, MAB treats each agent as an immutable stochastic arm, where one can query the arm (i.e.,  a model), observe a reward, and pay its cost, but cannot directly alter its internal parameters. We provide a specific implementation in Appendix C.1.
@@
In Appendix C.2, we propose several options for the construction of task banks and fundamental features for the initialization of the embedding. This dynamic and inter-dependent setting can be naturally modeled by temporal graph neural networks (TGNNs) [21, 22]. Each agent is a node, and each time-varying edge denotes an interaction (e.g., routing a sub-query or forwarding a partial answer) that arrives with a timestamp. TGNNs maintain a latent state for every node that is updated only when events involving that node occur, making them ideal for sparse, asynchronous communication patterns that dominate large-scale MAS.
@@
$\mathbf{z}_{i}^{t}$ $=\mathrm{Aggregate}\left(\{\mathbf{m}_{ij}^{t},j\in\mathcal{N}_{i}\}\right),\quad\mathbf{v}_{i}^{t}=\mathrm{Update}(\mathbf{z}_{i}^{t},\mathbf{v}_{i}^{t-1}),$ (4)
@@
[…]

### v3

@@ Abstract
Large Language Model-based Multi-Agent Systems (MASs) have emerged as a powerful paradigm for tackling complex tasks through collaborative intelligence. However, the topology of these systems–how agents in MASs should be configured, connected, and coordinated– remains largely unexplored. In this position paper, we call for a paradigm shift toward topology-aware MASs that explicitly model and dynamically optimize the structure of inter-agent interactions. We identify three fundamental components–agents, communication links, and overall topology–that collectively determine the system’s adaptability, efficiency, robustness, and fairness. To operationalize this vision, we introduce a systematic three-stage framework: 1) agent selection, 2) structure profiling, and 3) topology synthesis. This framework not only provides a principled foundation for designing MASs but also opens new research frontiers across language modeling, reinforcement learning, graph learning, and generative modeling to ultimately unleash their full potential in complex real-world applications. We conclude by outlining key challenges and opportunities in MASs evaluation. We hope our framework and perspectives offer critical new insights in the era of agentic AI.
Topological Structure Learning Should Be A Research Priority for LLM-Based Multi-Agent Systems
Jiaxi Yang1thanks: These authors contributed equally to this research. , Mengqi Zhang211footnotemark: 1 , Yiqiao Jin311footnotemark: 1 , Hao Chen4, Qingsong Wen5 Lu Lin1, Yi He2, Srijan Kumar3, Weijie Xu6, James Evans7, Jindong Wang2thanks: Corresponding to: Jindong Wang <jwang80@wm.edu> 1The Pennsylvania State University  2William & Mary  3Georgia Institute of Technology 4AMD  5Squirrel AI  6Amazon  7University of Chicago https://TopoAgent.github.io/
@@ 1 Introduction
Large Language Model (LLM) agents (Wu et al., 2023; Li et al., 2023; Liu et al., 2025) have demonstrated strong capabilities across diverse domains such as code generation (Hong et al., 2024; Islam et al., 2024; Qian et al., 2024), mathematical reasoning (Gou et al., 2023; Xie et al., 2024), and social simulation (Huang et al., 2024; Zhao et al., 2024). Building on the impressive capabilities of single-agents in general tasks, LLM-based Multi-Agent Systems (MASs) (Guo et al., 2024) have emerged as a powerful paradigm that coordinate multiple specialized agents–e.g. planners, retrievers, reasoners–to collaboratively solve complex, composite tasks. These systems assign distinct roles, such as planners, retrievers, and reasoners, to distinct agents, enabling human-like collaboration through a division of labor. Similar to a human organization, a MAS operates as a dynamically evolving ensemble that continually reconfigure to maintain a moving equilibrium. As agents join, depart, or evolve, the system adapts its collaboration structure to optimize for long-term performance under resource and latency constraints.
To enable inter-agent communication, coordination, and role specialization, MASs must support agents that are not only individually competent but also capable of collaborating through well-orchestrated interaction patterns. While individual agent competence has advanced rapidly, the organizational principle of how agents should be configured, connected, and coordinated–its topology–remains underexplored   (Qian et al., 2025). This reveals an urgent need to rethink MAS design: What topological structures of LLM-based MASs are best to accomplish an unseen, complex task?
Why Topology Matters?
In MASs, inter-agent communication determines how information flows, how tasks are decomposed, and how consensus are reached. As task complexity increases and the number of agents scale up, communication costs often rise quadratically Qian et al. (2025). Naïve scaling–adding more agents without adaptive structural design–often results in redundant communication, higher GPU memory and token costs, longer response latency, and suboptimal coordination (Qian et al., 2025; Zhang et al., 2024a; Wang et al., 2025).
(a) GPT‑3.5
(b) GPT‑4o
@@
Beyond efficiency concerns, topology also directly impacts task performance. As illustrated in Figure˜1, performance can vary by up to 10% across different topological structures. Furthermore, the choice of communication structure carries critical safety implications, whereby poorly designed topology may lead to misinformation propagation, biased reasoning, or the generation of harmful content by agents Yu et al. (2024).
Challenges
@@
First, the optimal topology is task-dependent. Universal topology design is difficult as structural preferences vary across domains–chain topologies are most effective for sequential workflows such as software development Zhang et al. (2024e); Hong et al. (2023), while simulation or reasoning tasks favor tree or star topology that support information aggregation Qian et al. (2025); Wu et al. (2023). These dependencies are implicit, context-sensitive, and hard to generalize across tasks. Although prior works explore dynamic topologies (Yue et al., 2025; Liu et al., 2024), they remain preliminary and still lack the ability to adaptively reconfigure topological structures in response to task goals or query contexts. Second, the design space of topological structures is inherently combinatorial. agent selection, communication edges, and role dependencies grow exponentially with scale, rendering exhaustive optimization infeasible. This makes efficient structure optimization computationally prohibitive as the system scales. Structure pruning techniques like edge pruning (Zhang et al., 2024a) and agent pruning (Wang et al., 2025; Liu et al., 2024) are often applied in a static, one-off manner, lacking the capacity to model the dynamic interdependencies and evolving interactions among agents. Third, evaluating topologies requires end-to-end execution under real task dynamics, which is computationally expensive and requires in-depth analytic modeling. Unlike in single-agent settings, the effectiveness of a given topology cannot be assessed in isolation. Its performance must be measured through full system execution within specific task contexts.
What Should be Learned and How?
We summarize three core components in MASs: agents, communication links, and the overall topology that governs how these agents interact. First, the candidate agent set should consist of capable and complementary specialists tailored to the task–balancing diversity of expertise with coordination efficiency. Second, communication links should be optimized to eliminate redundant messages and minimize communication overhead. Third, the topological structure must align with the task’s coordination pattern–whether hierarchical, modular, or distributed–to enable smooth information flow and balanced participation among agents.
Positions
In this paper, we advocate a structure-aware paradigm that treats topological design as a top priority LLM-based MASs. We propose a unified framework (Figure˜2) that decomposes topological structure optimization into three stages: agent selection (Section 3.2), structure profiling (Section 3.3), and topology synthesis (Section 3.4). This perspective reframes topology as a learnable, dynamic property–rather than a static configuration–opening new directions toward more adaptive, efficient, robust, and fair multi-agent systems.
Broader Impact
@@
Current agent-based LLM applications can potentially benefit from a topology-aware structure. For instance, in customer support automation, dynamic agent topologies can reduce response latency, minimize redundant information routing, and improve relevance in multi-turn dialogues for support systems involving billing, troubleshooting, and multilingual hand-offs. For collaborative code generation, task-specific topologies can support better coordination among agents with distinct roles (e.g., planner, coder, tester), which adapt as tasks evolve. Moreover, in scientific discovery, research assistants powered by LLMs can benefit from topologies that optimize how domain-specific agents (e.g., literature retrievers, experimental planners, hypothesis evaluators) interact, particularly when solving open-ended scientific queries. Given the dynamic and flexible nature of interaction topology, more applications will emerge and benefit.
@@
The organization of this paper is as follows: we first introduce the basics of LLM-based MASs and give the position statements in Section˜2. We then attempt to argue positions we stated in the three-step framework of topological structure learningin Section˜3. Subsequently, Section˜4 discusses the key challenges that explain why current MASs still fall short of realizing these topology-aware principles in practice. Finally, Section˜5 concludes the paper. Note that due to space limitations, more details of our proposed research directions are in the appendix.
@@
Figure 2: The proposed three-stage framework: agent selection, structure profiling, and topology synthesis, which grounds four key positions.
2 Positions
2.1 Definition of LLM-based Multi-Agent Systems
LLM-based single-agent systems excel on a wide range of tasks, but can struggle with long-horizon, tool-intensive, or multi-step problems. This motivates the development of MASs, where multiple LLM-based agents collaborate through natural language communication to constructively divide roles, specialize functions, and jointly solve tasks. To support such collaboration, we define an LLM-based MAS as a set of autonomous agents instantiated from LLMs, denoted as $\hat{\mathcal{A}}=\{a_{1},...,a_{i}\}\subseteq\mathcal{A}$. Each agent $a_{i}\in\mathcal{A}$ is assigned a specific role, such as planner, coder, or verifier, and interacts with other agents to collectively solve tasks. Appendix˜A shows our notations.
In practice, MASs adopt various structural topologies, such as chains, stars, trees, or densely connected graphs, each encoding different reasoning workflows and communication costs. For instance, chain structures enable step-by-step refinement (e.g., MetaGPT (Hong et al., 2023)), while tree structures support hierarchical planning (e.g., AutoGen Wu et al. (2023)). Star topologies simplify control via centralized agents (e.g., CAMEL (Li et al., 2023)), and fully connected graphs maximize information sharing (e.g., GPTSwarm (Zhuge et al., 2024)). Each structure presents distinct trade-offs in coordination efficiency, scalability, and reasoning capability, and is inherently better suited to different tasks depending on its specific communication patterns and collaborative requirements. Formally, we model the collaboration structure as a directed graph $\mathcal{G}=(\mathcal{A},\mathcal{E})$, where $\mathcal{E}\subseteq\mathcal{A}\times\mathcal{A}$ represents the communication links. At each round $t$, agent $a_{i}$ receives messages from its in-neighbors and generates an updated response $R_{i}^{(t)}=a_{i}(P_{\text{sys}},P_{\text{user}}^{(t)})$, where $P_{\textit{sys}}$, $P_{\text{user}}$ denote the system prompt and user prompt at round $t$ respectively. $P_{\text{user}}^{(t)}=P_{\text{user}}^{(0)}\cup\mathcal{O}_{i}^{(t-1)}$, where $\mathcal{O}_{i}^{(t-1)}$ denotes the aggregated messages from neighbors of agent $a_{i}$.
2.2 Position Statement
P 1.
(Resource Constraint) The topological structure in MASs directly determines the resource required for agent collaboration
Poorly designed topologies such as fully connected graphs lead to redundant communication, increasing VRAM usage, and increased latency, making collaboration costly, whereas centralized topologies such as star structures concentrate workloads on a minority of nodes, leading to computational bottlenecks.
P 2.
(Task performance) The topology of agents interactions serves as a critical factor in MASs that can either amplify or hinder task performance.
Different applications favor different agent structures for better performance. For instance, collaborative coding (e.g., ChatDev Qian et al. (2024) and AutoCodeRover Zhang et al. (2024d)) often uses chain structures for task decomposition, while world simulation (e.g., Minecraft Suwannik (2022) and ALFWorld Shridhar et al. (2021)) adopts a tree-based structure for better coordination.
P 3.
(Robustness) The robustness of MASs critically depends on their topological structure, which shapes the system’s tolerance to failures and adversarial perturbations.
Topologies with redundancy or modular design can sustain failures and resist misinformation. For example, debate structures introduce multiple verification paths that mitigate single-point failure. In contrast, fragile topologies amplify risks of collapse and error propagation, such as centralized star structures.
P 4.
(Fairness) The topology of agent interactions in MASs significantly influences the fairness of resource allocation and decision-making processes.
Balanced topologies enable diverse roles and personas (e.g., planners, critics, verifiers) to participate equitably, while centralized structures risk privileging certain agents and marginalizing others, leading to unfairness issues.
3 Position Arguments
To verify the four positions (P1–P4) we advocated, we ground our arguments in the three-step framework of topological structure learning: agent selection, structure profiling, and topology synthesis. At each stage, we discuss how topological structure considerations shape economic burden, task performance, robustness, and fairness. Beyond articulating these positions, we further outline potential research directions, demonstrating how each position can translate into actionable agendas for advancing MAS topology design.
3.1 Three-Stage Framework Overview
@@
To better substantiate the stated positions, we introduce a unified framework for learning the optimal topological structure of MASs and argue each position through the lens of its corresponding stage. The framework consists of three sequential stages: agent selection, structure profiling, and topology synthesis, as illustrated in Figure 2. Each stage has the potential to trigger new open problems and research directions. Specifically, ❶ Given a new task, a pool of candidate agents with diverse capabilities is instantiated and made available. The first step is to select a subset of agents $\hat{\mathcal{A}}\subseteq\mathcal{A}$ best suited to collaborate on the task, based on factors such as skill specialization, role diversity, and past performance. This determines “who” participates in the communication topology. ❷ For each task, the system identifies a macro structure as a communication pattern, which captures abstract interaction forms, such as chains, stars, or trees. These macro structures reflect coarse-grained collaboration strategies and serve as an initial prototype for the following topology synthesis. ❸ Based on steps ❶ and ❷, the final stage synthesizes a concrete and directed communication graph, which we reference as the micro structure. It specifies the exact flow of information among the selected agents under the given macro structure, and determines which agents exchange messages and in what direction.
@@ 3.2 Agent Selection
Selecting proper agents is a cost-utility optimization problem that must balance exploration (testing alternatives) and exploitation (using the best‑known option) (P1).
Users must decide which agents to choose as candidates–high-performance but expensive SaaS LLMs versus smaller local checkpoints–and how to allocate queries among agent. Under-exploration risks missing models with superior task performance (P2), while over-exploration and exhaustive trials induce redundancy and costs (P1).
@@
Existing agent selection schemes are fundamentally limited by three major challenges: Single-Agent Focus , Structural Myopia, and Scalability.
Users must balance when generalists and specialists fit into a particular task pipeline. We advocate a structure-aware, lifecycle-aware approach to agent selection that jointly optimizes multi-objective reward–task success, cost, latency, safety–while remaining adaptive to evolving agent availabilities and task distributions (P1, P2).
We propose two viable solutions, Static Selection and Dynamic Selection, depending on the specific scenarios. When task distribution, communication structure, evaluation metrics, costs, and the LLM catalog remain unchanged over the deployment horizon, e.g.,  nightly batch summarization with a locked‑in set of SaaS APIs, where the utility of an agent can be seen as stationary, the problem reduces to estimating and selecting an optimal subset based on the interaction structure, and thus Static Selection is preferred. Formally, assuming that each agent in the pool is associated with a cost $C(a_{i})\in\mathbb{R}_{\geq 0}$ (e.g., API credits, GPU runtime), for a downstream task $x$ under macro structure $M$, we can define a utility function $\mathcal{U}:(x,M,\hat{\mathcal{A}})\longmapsto[0,1]$ to evaluate the collective performance of a subset $\hat{\mathcal{A}}\subseteq\mathcal{A}$ on task $x$. This utility can reflect metrics such as answer accuracy, semantic relevance, or user engagement (e.g., click-through rate). The goal of static selection to find $k$ agents to balance performance-cost trade-offs:
[…]
$\quad-\sum_{a_{i}\in\hat{\mathcal{A}}}\gamma(a_{i})C(a_{i})\Bigr],$
@@
The multiplier $\gamma$ in Equation 3.2 can be treated as a stochastic, context-dependent signal, enabling adaptive selection policies that dynamically balance performance and economy as conditions drift.
Multi‑armed bandits (MAB) and their combinatorial variants offer a principled framework to efficiently explore promising agents and exploit high-utility subsets. Unlike reinforcement‑learning (RL) training loops that continuously update a policy’s parameters, MAB treats each agent as an immutable stochastic arm, where one can query the arm (i.e.,  a model), observe a reward, and pay its cost, but cannot directly alter its internal parameters. We provide a specific implementation in Appendix C.1.
@@
In Appendix C.2, we propose several options for the construction of task banks and fundamental features for the initialization of the embedding. This dynamic and inter-dependent setting can be naturally modeled by temporal graph neural networks (TGNNs) (Rossi et al., 2020; Kumar et al., 2019). Each agent is a node, and each time-varying edge denotes an interaction (e.g., routing a sub‑query or forwarding a partial answer) that arrives with a timestamp. TGNNs maintain a latent state for every node that is updated only when events involving that node occur, making them ideal for sparse, asynchronous communication patterns that dominate large-scale MAS.
@@
$\mathbf{z}_{i}^{t}$ $=\mathrm{Agg}\left(\{\mathbf{m}_{ij}^{t},j\in\mathcal{N}_{i}\}\right),$ (3)
$\mathbf{v}_{i}^{t}$ $=\mathrm{Update}(\mathbf{z}_{i}^{t},\mathbf{v}_{i}^{t-1}),$ (4)
@@
where $\|$ is concatenation. The subset $\hat{\mathcal{A}}$ is scored via automatic or LLM-based evaluation as a function of its utility. Both the TGNN and the read-out module are trained jointly using a pairwise ranking loss that promotes higher scores for subsets yielding greater utility. The model learns to predict both the subset utility and each node’s marginal contribution.
A fairness-aware read-out module can be introduced to prevent systematic bias toward resource-rich or proprietary agents. Without such correction, the model might consistently prefer high-capacity agents (e.g., GPT-4-class LLMs) that dominate short-term utility estimates, while overlooking smaller open-source agents that provide complementary reasoning diversity or cost efficiency (P4).
Along this direction, the readout can incorporate group-level regularization or exposure balancing—e.g., ensuring that subsets include a mix of agent types (proprietary vs. open-source), or applying a mild penalty when all selected agents belong to a single high-cost group.